# An operational approach to classifying measurement incompatibility

### v1

@@ III.1 Coarse-graining of outcomes
$|\psi_{2}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega^{2}|1\rangle+\omega|2\rangle),$
It is observed that for $\nu>0.683$, the two measurements are incompatible. For finding ”$2$-incompatible w.r.t coarse-graining”, we consider all possible coarse-graining of outcomes of the two measurements such that after coarse-graining, the new set of measurements have at least two outcomes each. We find that these measurements are incompatible w.r.t all possible coarse-graining for $\nu>0.711$. Note that for $0.683<\nu\leq 0.711$, the incompatible measurements become compatible under certain coarse-graining of outcomes (see TABLE 1).
Dimension $4$. For checking incompatibility w.rt coarse-graining in $\mathbb{C}^{4}$, the same procedure is repeated for two POVM measurements with $4$ outcomes each. The corresponding measurements are $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\},i,j\in[4]$.
$P^{\prime}_{i}$ $=$ $\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{4},$
$Q^{\prime}_{j}$ $=$ $\nu|\psi^{\prime}_{j}\rangle\langle\psi^{\prime}_{j}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (24)
@@
$|\psi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+|2\rangle-|3\rangle),$ (25)
for $\nu>0.666$, these measurements are incompatible. We find that for $\nu>0.675$, the measurements are incompatible under all possible non-trivial coarse-graining that gives rise to at least $3$ outcomes. The measurements are fully incompatible for $\nu>0.72$. The results are summarized in TABLE 1.
@@
3-incompatible 0.683 0.675
Table 1: Critical values of robustness for MUBs w.r.t coarse-graining in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
III.2 Convex mixing of measurements
Dimension $3$. For checking incompatibility w.r.t convex mixtures (CM), a minimum of three measurements are needed. The two measurements are $\{P_{i}\}$ and $\{Q_{j}\}$, as given in Eqs.(23), The third measurement is $\{R_{k}\},k\in[3]$ where
@@ III.2 Convex mixing of measurements
$\begin{split}&|\phi_{0}\rangle=\frac{\omega|0\rangle+|1\rangle+|2\rangle}{\sqrt{3}},|\phi_{1}\rangle=\frac{|0\rangle+\omega|1\rangle+|2\rangle}{\sqrt{3}},\\ &|\phi_{2}\rangle=\frac{|0\rangle+|1\rangle+\omega|2\rangle}{\sqrt{3}},\end{split}$ (27)
$\omega$ being the cube roots of unity.
The results are summarized in TABLE 2.
Dimension $4$. Consider the measurements $\{P^{\prime}_{i}\},\{Q^{\prime}_{j}\}$ as defined in Eqs.(III.1) and(III.1) and another measurement $\{R^{\prime}_{k}\}$ where $i,j,k\in[4]$.
$\{R^{\prime}_{k}\}=\nu|\phi^{\prime}_{k}\rangle\langle\phi^{\prime}_{k}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (28)
@@
The results are summarized in TABLE 2.
@@
3-incompatible 0.537 0.692
Table 2: Critical values of robustness for MUBs w.r.t convex-mixing in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$.
IV Operational witnesses of incompatibility
IV.1 Coarse-graining of outcomes
As we mentioned earlier, measurements that are fully incompatible w.r.t coarse-graining show stronger incompatibility compared to the measurements that are not fully incompatible w.r.t coarse-graining. To operationally certify whether a set of measurements are fully incompatible w.r.t coarse-graining is important from the perspective of determing the practical utility of such a set for revealing phenomena such as Bell inequality, steering and contextuality Fine (1982a); Wolf et al. (2009); Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015); Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), as well as for checking the proficiency of such a set for information processing tasks Carmeli et al. (2020); Saha et al. (2023). Operational certification means that we need to infer the incompatibility of the measurements from the input-output statistics of the measurement device without knowing its internal functioning. Below we study the two classes of operational certification separately
@@ IV.1 Coarse-graining of outcomes
Here one requires neither any prior knowledge of the internal functioning of the measurement device, nor any idea of the dimension of the system on which the measurements act. In Wolf et al. (2009), it has been proven that any two binary-outcome incompatible measurements violate the Bell-CHSH inequality. We can use this result to witness incompatibility for any two 3-outcome projective measurements before and after coarse-graining, from the input-output statistics.
Semi-Device-independent witness
Here, although we do not have any prior knowledge of the internal functioning of the measurement device, we do know the dimension of the system on which the measurements act. Any two incompatible rank-one projective measurements provide an advantage in $(2,d,d)-$RAC Saha et al. (2020). If any two POVMs each with $\overline{d}$ outcomes acting on $\mathbb{C}^{d}$ are jointly measurable, the average maximum success probability in $(2,\overline{d},d)$ RAC is bounded by Saha et al. (2023)
$P(2,\overline{d},d)\leq\frac{1}{2}(1+\frac{d}{\overline{d}^{2}})=P_{CB}(2,\overline{d},d),$ (30)
where $P_{CB}(2,\overline{d},d)$ is the maximum average success probability obtained in a $(2,\overline{d},d)$ RAC using two compatible measurements. For full incompatibility of two 3-outcome rank-one projective measurements, we prove the following:
Theorem 3.
Two rank-one projective measurements, $M=\Big{\{}|\phi_{0}\rangle,|\phi_{1}\rangle,|\phi_{2}\rangle\Big{\}}$ and $N=\Big{\{}|\psi_{0}\rangle,|\psi_{1}\rangle,|\psi_{2}\rangle\Big{\}}$, can be witnessed to be fully incompatible w.r.t. coarse-graining via RAC iff $20<|\langle\phi_{i}|\psi_{j}\rangle|<\frac{4}{5},\quad\forall\,\,i,j=0,1,$ 2 .
@@
$M=\Big{\{}|0\rangle\langle 0|,|1\rangle\langle 1|,|2\rangle\langle 2|\Big{\}}\equiv\Big{\{}|\phi_{i}\rangle\langle\phi_{i}|\Big{\}},$
and
$N=\Big{\{}|\psi_{1}\rangle\langle\psi_{1}|,|\psi_{2}\rangle\langle\psi_{2}|,|\psi_{3}\rangle\langle\psi_{3}|\Big{\}}\equiv\Big{\{}|\psi_{j}\rangle\langle\psi_{j}|\Big{\}},$ (31)
Let us first coarse grain the second and third outcomes. So, the new measurement pair becomes:
The new compatibility bound in this case is,
@@
and the bound is tight. $M^{(2,3)}\,\,\text{and}\,\,N^{(2,3)}$ will give advantage in the (2,2,3) RAC if:
$\sum_{x,y=0}^{1}$ $||M^{(2,3)}(x)+N^{(2,3)}(y)||>7,\,\,\text{which implies}$
$||A||$ $+$ $||B||+||C||+||D||>7$ (33)
$\text{with}\,\,A$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(0)$
$=$ $|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
Now, $|\psi_{1}\rangle$ can be written as:
$\langle 0|u\rangle=0\,\,\text{and}\,\,|\langle 0|\psi_{1}\rangle|^{2}+|\langle u|\psi_{1}\rangle|^{2}=1.$
$A$ can be written as a $(2\times 2)$ matrix in the basis $\{|0\rangle,|u\rangle\},$
$A=\begin{pmatrix}1+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (37)
So, the maximum eigenvalue of $A\,\,\text{is}\,\,1+|\langle 0|\psi_{1}\rangle|.$ Thus,
$||A||=1+|\langle 0|\psi_{1}\rangle|.$
By Eq.(IV.1) we have
$B$ $=$ $M^{(2,3)}(0)+N^{(2,3)}(1)$ (38)
$=$ $|0\rangle\langle 0|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|$
$=$ $\mathbbm{1}+|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|.$
Similarly, $B$ can be expressed in a block diagonal matrix in the ortho-normal basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$B=\begin{pmatrix}\Gamma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Gamma=\begin{pmatrix}2+|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (43)
and
$||B||=2+|\langle 0|\psi_{1}\rangle|.$
Also, from Eq.(IV.1)
$C$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(0)$ (44)
$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{1}\rangle\langle\psi_{1}|\,\,$
$=$ $\mathbbm{1}-|0\rangle\langle 0|+|\psi_{1}\rangle\langle\psi_{1}|.$
Similarly, $C$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$C=\begin{pmatrix}\Sigma&0\\ 0&1\end{pmatrix}\,\,\text{where}\,\,\Sigma=\begin{pmatrix}|\langle 0|\psi_{1}\rangle|^{2}&\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ \langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&1+|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (49)
and
$||C||=1+\sqrt{1-|\langle 0|\psi_{1}\rangle|^{2}}.$
Similarly, from Eq.(IV.1)
$D$ $=$ $M^{(2,3)}(1)+N^{(2,3)}(1)$ (50)
$=$ $\,\,|1\rangle\langle 1|+|2\rangle\langle 2|+|\psi_{2}\rangle\langle\psi_{2}|+|\psi_{3}\rangle\langle\psi_{3}|\,\,$
$=$ $2\mathbbm{1}-|0\rangle\langle 0|-|\psi_{1}\rangle\langle\psi_{1}|,$
and $D$ can be expressed in a block diagonal matrix in the basis $\{|0\rangle,|u\rangle,|v\rangle\},$
$D=\begin{pmatrix}\Xi&0\\ 0&2\end{pmatrix}\,\,\text{where}\,\,\Xi=\begin{pmatrix}1-|\langle 0|\psi_{1}\rangle|^{2}&-\langle 0|\psi_{1}\rangle\langle\psi_{1}|u\rangle\\ -\langle u|\psi_{1}\rangle\langle\psi_{1}|0\rangle&2-|\langle u|\psi_{1}\rangle|^{2}\end{pmatrix},$ (55)
and $||D||=2.$
[…]
Device-independent witness of full-incompatibility of two measurements w.r.t coarse-graining from a single experiment
We now study the following problem: given two measurements of outcomes $d_{1}$ and $d_{2}$ respectively, how can one witness the full-incompatibility w.r.t coarse-graining in a device-independent way from a single experiment? Here ”single experiment” implies the Bell test which gives a Bell-inequality violation for the two measurements of outcomes $d_{1}$ and $d_{2}$ respectively as chosen as inputs of Alice and some measurement choices for Bob and a shared entangled state. As we know if two measurements are $2-$ incompatible w.r.t coarse-graining then they are fully-incompatible w.r.t coarse-graining. Also, we use the result of Wolf et al. (2009), that any two binary outcome incompatible measurements violate the Bell-CHSH inequality. Thus, from the original measurement statistics we do all possible coarse-grainings to make it two outcome measurements and check the Bell-CHSH inequalities. If, all possible coarse-grained two-outcome measurements violate the Bell-CHSH inequalities then the measurements are fully-incompatible w.r.t coarse-graining.
To understand how this technique works, we consider a well-known example of two three-outcome rank-one projective measurements in $\mathbb{C}^{3}$ which give maximum violation of the CGLMP inequality Collins et al. (2002), a well-studied Zohren and Gill (2008); Das et al. (2019); Karczewski et al. (2022) Bell-type inequality. The projective measurements of Alice are: $A_{a}\equiv\{|\xi\rangle_{A,a}\},$ where
@@
The incompatibility of the measurements of Eq.(56) before coarse-graining is guaranteed by the CGLMP-inequality violation. Now we investigate the incompatibility status of the measurements of Alice after all possible coarse-graining from the measurement statistics obtained in the CGLMP experiment. Alice’s measurements now have two outcomes each and it is a (2,2,2,3) scenario, i.e., two inputs for Alice and two inputs for Bob and each input of Alice has two outcomes and each input of Bob has three outcomes. For this scenario, Collins and Gisin have shown that there are a total of 72 CH-facet inequalities Collins and Gisin (2004). The CH-inequalities Fine (1982b) are of the form:
$-1\leq$ $P(00|A_{1},B_{1})+P(00|A_{1},B_{2})+P(00|A_{2},B_{2})$ (58)
$-P(00|A_{2},B_{1})-P(0|A_{1})-P(0|B_{2})\leq 0.$
@@
$P(\overline{0}|B_{i})=P(0|B_{i})+P(1|B_{i}),\,\,i,j\in\{1,2\}.$ (61)
@@
Similarly, there are other 8 possible clubbings for Bob’s measurement outcomes and for each clubbing, we have facet inequalities similar to Eq.(59). One can check that when there are the same coarse-graining of outcomes for both measurement inputs of Alice, we get a CH-inequality violation. Thus, under these coarse-grainings, the two measurements of Alice remain incompatible. When there are different coarse-grainings, for some cases we get CH-violation, but for other cases, we do not get CH-violation. CH-violation under a particular coarse-graining signifies that the measurements are incompatible. However, when there is no CH-violation we can not conclude anything regarding the incompatibility status. This is depicted in TABLE   3.
$A_{1}$ $A_{2}$ Incompatible
(0,1) (0,1) yes
(0,1) (1,2) ?
(0,1) (0,2) yes
(1,2) (0,1) yes
(1,2) (1,2) yes
(1,2) (0,2) ?
(0,2) (0,1) ?
(0,2) (1,2) yes
(0,2) (0,2) yes

### v2

@@ IV.1 Robustness of incompatibility under coarse-graining of outcomes
$|\psi_{2}\rangle=\frac{1}{\sqrt{3}}(|0\rangle+\omega^{2}|1\rangle+\omega|2\rangle),$ (28)
$\omega$ being the cube roots of unity. It can be easily checked by SDP that the measurements in Eqs.(27) and (28) are incompatible (or equivalently, $3$-incompatible using the definition 2 ) for robustness parameter $(\nu)>0.683$. Now, any non-trivial coarse-graining reduces the measurement outcomes to two, so one can check the incompatibility after all possible choices of such non-trivial coarse-grainings. The measurements are $2$-incompatible (by the definition 2) for $\nu>0.711$. For $0.683<\nu\leq 0.711$, the measurements are $3$-incompatible, but not $2$-incompatible. This represents the case where the initial measurements, although taken to be incompatible, become compatible after a non-trivial coarse-graining. However, for $\nu>0.711$, the measurements are $2$-incompatible (and hence $3$-incompatible), suggesting that they remain incompatible after all non-trivial coarse-grainings. These findings suggest that $2$-incompatible measurements exhibit a stronger form of incompatibility compared to the $3$-incompatible measurements (hence supporting Observation 1), and hence are potential candidates for any resource-theoretic applications that harness incompatibility. Note that for three outcome measurements, any non-trivial coarse-graining will reduce the outcomes to two. So, the number of outcomes is always $\leq 3$ (where $3$ corresponds to the case where there is no coarse-graining). So, $4$-incompatibility is not applicable (N.A.) here.
Dimension $4$. For checking incompatibility w.r.t coarse-graining in $\mathbb{C}^{4}$, the same procedure is repeated taking two POVM measurements with four outcomes each. The corresponding measurements are $\{M^{\prime}_{i}\},\{N^{\prime}_{j}\},i,j\in\{1,2,3,4\}$:
$M^{\prime}_{i}$ $=$ $\nu|i\rangle\langle i|+(1-\nu)\frac{\mathbbm{1}}{4},$
$N^{\prime}_{j}$ $=$ $\nu|\psi^{\prime}_{j}\rangle\langle\psi^{\prime}_{j}|+(1-\nu)\frac{\mathbbm{1}}{4},$ (29)
@@
$|\psi^{\prime}_{3}\rangle$ $=$ $\frac{1}{2}(|0\rangle-|1\rangle+|2\rangle-|3\rangle).$ (30)
These measurements in Eqs.(IV.1) and (IV.1) are incompatible (or equivalently $4$-incompatible, using the definition 2) for $\nu>0.666$. In this case, any non-trivial coarse-graining will reduce the measurement outcomes to three or two, and hence we shall have $3$-incompatiblity and $2$-incompatiblity respectively. For $0.666<\nu\leq 0.675$, the incompatible measurements become compatible after all possible coarse-grainings that reduce the outcomes to at least three. Further reducing the number of outcomes by coarse-graining gives a tighter bound for the critical value of the robustness parameter above which the measurements are incompatible. Hence, for $0.675<\nu\leq 0.720$, the measurements are $3$-incompatible but become compatible if one further reduces the number of outcomes by coarse-graining. These measurements exhibit the strongest form of incompatibility for $\nu>0.720$ (as also discussed in Observation 1), i.e., they remain incompatible after all possible non-trivial coarse-grainings. The results for coarse-graining of measurement outcomes in dimensions $3$ and $4$ are summarized in TABLE 1.
@@
Table 1: Critical values of robustness for MUBs w.r.t. coarse-graining in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$, above which the measurements are incompatible.
IV.2 Robustness of incompatibility under disjoint-convex-mixing of measurements
Dimension $3$. For checking incompatibility w.r.t. disjoint-convex-mixing (CM), a minimum of three measurements are needed. The two measurements are $\{M_{i}\}$ and $\{N_{j}\}$, as given in Eqs.(27) and Eq.(28). The third measurement is $\{R_{k}\},k\in\{1,2,3\}$ where
@@ IV.2 Robustness of incompatibility under disjoint-convex-mixing of measurements
$\begin{split}&|\phi_{0}\rangle=\frac{\omega|0\rangle+|1\rangle+|2\rangle}{\sqrt{3}},\ |\phi_{1}\rangle=\frac{|0\rangle+\omega|1\rangle+|2\rangle}{\sqrt{3}},\\ &|\phi_{2}\rangle=\frac{|0\rangle+|1\rangle+\omega|2\rangle}{\sqrt{3}}.\end{split}$ (32)
These three measurements are incompatible (equivalently $3$-incompatible, by the definition 4) for $\nu>0.537$. From these three incompatible measurements, one can effectively reduce the number of measurements to two, by taking a new measurement as the disjoint-convex-mixing of any two measurements from the initial set of three measurements (this can be done in $3$ ways), while keeping the third measurement same. This process of disjoint-convex-mixing may reduce the degree of incompatibility, as discussed before and evidenced for $0.537<\nu\leq 0.764$. This corresponds to the case where the two measurements become compatible after considering all possible disjoint-convex-mixing across all partitions. For $\nu>0.764$, the measurements are incompatible, suggesting that this is the critical value above which, the measurements are robust against disjoint-convex-mixing.
Dimension $4$. Consider the measurements $\{M^{\prime}_{i}\},\{N^{\prime}_{j}\}$ as defined in Eqs. (IV.1) and (IV.1) and another measurement $\{R^{\prime}_{k}\}$ where $i,j,k\in\{1,2,3,4\}$,
$R^{\prime}_{k}=\nu|\phi^{\prime}_{k}\rangle\langle\phi^{\prime}_{k}|+(1-\nu)\frac{\mathbbm{1}}{4}$ (33)
@@
These three measurements in dimension $4$ are incompatible for $\nu>0.692$. As discussed before, disjoint-convex-mixing makes them compatible and they continue to remain so till $\nu=0.705$. So, for $0.692<\nu\leq 0.705$, the three measurements are incompatible, but disjoint-convex-mixing of any two of them makes them compatible. For $\nu>0.705$, the measurements are incompatible and robust against disjoint-convex-mixing across any partition that effectively brings down the number of measurements to two. The results for disjoint-convex-mixing of measurements in dimensions $3$ and $4$ are summarized in TABLE 2.
@@
Table 2: Critical values of robustness for MUBs w.r.t. disjoint-convex-mixing in $\mathbb{C}^{3}$ and $\mathbb{C}^{4}$, above which the measurements are incompatible.
V Operational witnesses of incompatibility
In this section, we explore how different levels of incompatibility can be witnessed through the input-output statistics derived from characterized devices.
V.1 Coarse-graining of outcomes
As we mentioned earlier, measurements that are fully incompatible w.r.t. coarse-graining show stronger incompatibility compared to the measurements that are not fully incompatible w.r.t. coarse-graining. To operationally certify whether a set of measurements is fully incompatible w.r.t. coarse-graining is important from the perspective of determining the practical utility of such a set for revealing phenomena such as Bell inequality, steering and contextuality Fine (1982a); Wolf et al. (2009a); Uola et al. (2014); Quintino et al. (2014); Pramanik et al. (2014); Uola et al. (2015); Chowdhury et al. (2015); Liang et al. (2011); Xu and Cabello (2019); Gupta et al. (2023), as well as for checking the proficiency of such a set for information processing tasks Carmeli et al. (2020); Saha et al. (2023). Certification means that we need to infer the incompatibility of the measurements from the input-output statistics of the measurement device without knowing its internal functioning. Below we study the two classes of certification separately.
@@ V.1 Coarse-graining of outcomes
Violation of Bell inequality provides device-independent witness for incompatible measurements. Here, one neither requires any prior knowledge of the internal functioning of the measurement device nor any idea of the dimension of the system on which the measurements act. However, not all incompatible measurements yield violations of Bell inequalities Bene and Vértesi (2018); Hirsch et al. (2018). Nonetheless, we have the following results.
Observation 4.
Full-incompatibility w.r.t. coarse-graining of any two measurements can always be witnessed in a device-independent way.
@@
Device-independent witness from a single Bell experiment
In the above-mentioned approach, different Bell experiments are probed for different coarse-graining. It is interesting to explore whether different levels of incompatibility w.r.t. coarse-graining can be witnessed universally through a single Bell experiment in a device-independent manner. In this case, the objective is to consider different coarse-grained statistics of a Bell experiment and check whether those statistics have local explanations or not.
To understand how this technique works, we consider a well-known example of two three-outcome rank-one projective measurements in $\mathbb{C}^{3}$ which give maximum violation of the Collins-Gisin-Linden-Massar-Popescu (CGLMP) inequality Collins et al. (2002), a well-studied Zohren and Gill (2008); Das et al. (2019); Karczewski et al. (2022) Bell-type inequality. The projective measurements of Alice are: $A_{a}\equiv\{|\xi\rangle_{A,a}\},$ where
@@
Specifically, here we investigate the different layers of incompatibility of the measurements given in (35) under coarse-graining. The incompatibility of the measurements of Eq.(35) before coarse-graining is guaranteed by the CGLMP-inequality violation. Now we give a theoretical analysis to investigate the incompatibility status of the measurements of Alice given in (35) after all possible coarse-grainings. The results are summarized in TABLE   3.
Theoretical explanation.— Out of three outcomes if any of the two outcomes are coarse-grained it becomes a (2,2,2,3) scenario, i.e., two inputs for Alice and two inputs for Bob and each input of Alice has two outcomes and each input of Bob has three outcomes. For this scenario, Collins and Gisin have shown that there are a total of 72 CH-facet inequalities Collins and Gisin (2004). The CH-inequalities Fine (1982b) are of the form:
$-1\leq S\leq 1,$ (37)
with CH-functional
$S=$ $P(00|A_{1},B_{1})+P(00|A_{1},B_{2})+P(00|A_{2},B_{2})$ (38)
$-P(00|A_{2},B_{1})-P(0|A_{1})-P(0|B_{2}).$
@@
$P(\overline{0}|B_{j})=P(0|B_{j})+P(1|B_{j}),\,\,i,j\in\{1,2\}.$ (41)
Similarly, there are eight other possible clubbings for Bob’s measurement outcomes and for each clubbing, we have facet inequalities similar to Eq.(39).
@@
One can check that when there are the same coarse-graining of outcomes for both measurement inputs of Alice, we get a CH-inequality violation. Thus, under these coarse-grainings, the two measurements of Alice remain incompatible. When there are different coarse-grainings, for some cases we get CH-violation, but for other cases, we do not get CH-violation. CH-violation under a particular coarse-graining signifies that the measurements are incompatible. However, when there is no CH-violation, we can not conclude anything regarding the incompatibility status. This is depicted in TABLE   3.
Possible
CG
choices $\Delta S_{T}$ $\Delta S_{E}$ Incompatible
of
$A_{1}$ $A_{2}$
(0,1) (0,1) 0.126 0.122 yes
(0,1) (1,2) 0 0 ?
(0,1) (0,2) 0.126 0.122 yes
(1,2) (0,1) 0.126 0.122 yes
(1,2) (1,2) 0.126 0.122 yes
(1,2) (0,2) 0 0 ?
(0,2) (0,1) 0 0 ?
(0,2) (1,2) 0.126 0.122 yes
(0,2) (0,2) 0.126 0.122 yes